OpenStudy (gabrielah96):

4arcsinx-4pi= 2arcsin5pi. find exact value. can someone show me all the steps?

OpenStudy (anonymous):

not sure i can read it \[4\sin^{-1}(x)-4\pi=2\sin^{-1}(5\pi)\]?

OpenStudy (anonymous):

that can't be right, must be something else

OpenStudy (gabrielah96):

the first half is correct the 2nd half is 2arcsin x -5pi

OpenStudy (gabrielah96):

I'm sorry i typed it wrong

OpenStudy (anonymous):

isolate the arcisne first like solving \[4x-4\pi=2x-5\pi\] for \(x\)

OpenStudy (anonymous):

clear or no?

OpenStudy (gabrielah96):

yes I'm trying it

OpenStudy (gabrielah96):

i got -pi/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you have \[\sin^{-1}(x)=-\frac{\pi}{2}\]

OpenStudy (anonymous):

making \[x=\sin(-\frac{\pi}{2})\]

OpenStudy (gabrielah96):

that wasn't one of the answer choices

OpenStudy (gabrielah96):

would it be -1/2?

OpenStudy (anonymous):

noi it is not -1/2

OpenStudy (anonymous):

look on the unit circle to find \[\sin(-\frac{\pi}{2})\]

OpenStudy (gabrielah96):

oh okay i see it now so it would be -1

OpenStudy (anonymous):

yes

OpenStudy (gabrielah96):

thank you!